# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v5

@@ Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we developed CataractBot [^1] Our code is publicly available at https://github.com/microsoft/BYOeB.. CataractBot answers cataract surgery related questions instantly using an LLM to query a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 49 patients and attendants, 4 doctors, and 2 patient coordinators, CataractBot demonstrated potential, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
@@
$c$s: Human-centered computing Interaction designccs: Applied computing Health care information systemsccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
To address patients’ information needs, particularly during surgical treatments, we propose a chatbot solution integrated with medical professionals to provide accessible, reliable and comprehensive responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of smartphones and instant messaging services such as WhatsApp, this chatbot serves as a 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified this experts-in-the-loop chatbot approach by developing CataractBot, which answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (Lewis et al., 2020) with an LLM over a custom knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features to enhance adoption and sustained use. Capitalizing on the widespread adoption of smartphones, we chose a mobile-first design paradigm, deploying the chatbot entirely on the ubiquitous WhatsApp platform. To cater to diverse hospital visitors with varying literacy levels and technological proficiency, we made CataractBot multilingual (supporting five languages, including English, Hindi, and Kannada) and multimodal (accepting both speech and text inputs). We use an LLM to process complex and ill-formed queries, generating instant responses from the curated knowledge base. Crucially, CataractBot incorporates a system where the LLM’s responses are asynchronously verified by experts and corrected if needed. Ophthalmologists verify answers to medical queries and patient coordinators to logistical ones. These expert-provided edits are used to update the knowledge base, to minimize future expert intervention.
@@
Our main contributions are: (1) The design and development of a novel LLM-powered, experts-in-the-loop, multilingual, multimodal chatbot, created in collaboration with doctors and hospital staff to assist cataract surgery patients with their information needs. (2) The deployment of our bot in a real-world multi-stakeholder setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2.2. LLM Chatbots in Healthcare
We chose cataract surgery–one of the world’s most common procedures (McGhee et al., 2020)–as the use case for developing and testing our system. This decision was motivated by our longstanding collaboration with Sankara Eye Hospital, an eye hospital in India. Also, cataract surgery involves well-defined protocols and common patient queries, making it an ideal starting point for testing the bot’s ability to deliver accurate, standardized information from a curated knowledge base. However, our solution is versatile, and we believe it can effectively support a range of health conditions, including mental health and chronic care. Given our specific focus on cataract surgery, we examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
@@ 3.2. Design Requirements
Usability. After cataract surgery, patients can reference their discharge summaries for any queries, but this is challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role family members play in monitoring the patient’s health. Coupled with doctors’ input (“Patients ask hardly 10% of the question… attendants are the ones who ask.”), this underscored the necessity for the bot to be usable by both the patients and their attendants. To cater to diverse backgrounds, we made the chatbot multilingual and multimodal. We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022) and therefore minimal learning curve and quick onboarding without additional installations, despite the richer feature set of other platforms like Telegram or the development flexibility of a custom chatbot. Considering doctors’ demanding schedules, often involving consulting $\sim$50 patients and performing $\sim$10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s knowledge base, minimizing similar edits in future.
@@ 4. CataractBot System Design
Figure 1. CataractBot provides an initial response to the patient/attendant by querying the knowledge base. The doctor (or coordinator, if the question is logistical) verifies and corrects this response, and the patient/attendant is notified.:
@@ 4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk (e.g., greetings or expressions of gratitude), or ‘other’, and provides a response in real-time. For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response. This custom knowledge base comprises of various cataract surgery and hospital-specific documents curated by hospital staff (doctors, patient coordinators, and members of Quality Control and the Patient Safety team). These documents, comprising approximately 30 pages in total, include the Consent Form, Standard Operating Procedures, FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. Some documents, like Post-Operative Guidelines, were included in patients’ discharge summaries. Others, such as SOPs and FAQs, existed in the hospital ecosystem but were not directly accessible to patients. The generated response includes a question mark icon [Uncaptioned image] as a ‘reaction’ indicating the unverified status (Figure 4E in Appendix A.1.2). In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response. For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
@@ 4.2. Implementation Details
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and knowledge base expert) get logged in a cloud storage for further analysis.
Figure 2. Overview of CataractBot System:
@@ 5.1. Procedure for Patients and Attendants
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking $\sim$45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
@@ 6.1.1. Reasons for (Not) Using CataractBot
Figure 3. A. (Left) Messages sent per day pre- and post-surgery. B. (Top right) Messages sent before surgery, on the day of surgery, and after surgery. C. (Bottom right) Message input modality.:
@@ 7. Discussion
Below, we identify design implications, raising critical insights and open questions that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale across domains.
7.1. Design Implications
7.1.1. Generalization and Scaling
@@
Our open-sourced framework can be adapted beyond cataract surgery, to support more complex, atypical, or long-term treatments–e.g., glaucoma surgery, cancer treatment, or pregnancy–where easy access to verified information would be valuable for patients. The applicability can be extended further, to fields beyond the medical domain, such as law, finance and education. In these scenarios, end-users (e.g., defendants, taxpayers, or students) could receive synchronous responses from an LLM using a custom knowledge base, and subsequent asynchronous verified responses from experts (e.g., lawyers, accountants, or teachers). Expert-mediated bots can even be used to upskill less-trained workforce members. For instance, these bots could serve as valuable learning resources for community health workers, who typically have minimal medical training (Yadav et al., 2019).
7.1.2. Expert-in-the-loop Approaches
@@
Doctors are a scare resource in healthcare ecosystems. Our chatbot system can help scale their impact, by reducing the time spent answering routine questions at length during consultations and calls, thus freeing up time for additional patient care. However, as doctors had to verify responses during deployment, CataractBot also added to their workload. To minimize this, one approach is to deliver pre-verified answers to specific questions that the expert has repeatedly marked as correct. Given the potential for LLM hallucinations (Denecke et al., 2024), cached responses must be carefully matched to similar patient demographics and question phrasing before reuse. Another approach is to use a second LLM for quality control while generating each response, as GPT-4’s evaluations of AI-generated responses align well with human expert assessments (Gumma et al., 2024). This maker-checker verification module would enable tracking of the bot’s response quality, especially as the knowledge base expands. The module could also evaluate the confidence level in the generated response, or the perceived urgency of the question, which could be used to prioritize sending nudges to experts for verification.
7.1.3. Multimodality and Personalization
@@
Our system faced challenges in accurately interpreting questions from voice messages due to limitations in transcription and translation technologies. The same inaccurate transcriptions were shared with doctors leading to confusion. Such issues would not arise in deploying a voice-based input modality to a predominantly English-speaking population or a tier-1 language setting in the Global North, as seen in prior work (Yang et al., 2024). Additionally, non-verbal cues in audio messages–such as signs of anxiety, impatience, or the presence of others–were lost in textual transcriptions. This hindered experts’ ability to deliver appropriate care (also noted in (Li et al., 2023)). To address this, we suggest making the original audio clips available to experts on-demand. Beyond audio, considering the visual capabilities of recent LLM models like GPT-4V, such bots could be enhanced to process image-based inputs, aligning with recommendations in prior studies (Jain et al., 2018). For example, a patient could upload a photo of a specific medicine to ask questions about it, instead of providing a written description. While this could improve the user experience for information seekers, it may increase the workload for experts. Further, prior research (Hao et al., 2024) cautions that although LLMs are proficient in text processing, their ability to handle other forms of media, such as image and video, remains less reliable.
@@
Moving on to personalization, while tailored responses can boost adoption and utility by providing more relevant information, they also raise privacy concerns, potentially reducing usage among certain users. This issue may be more pronounced in the developing world, where there is more awareness around privacy. Additionally, experts may face increased workloads as they reference user-specific information for verification, since many personalized edits may not be suitable for inclusion in the generic knowledge base. Low (tech-)literate users might require assistance in making informed decisions about the use of their data (Okolo et al., 2021). A binary approach of either sharing all records or none may not address individual preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access. This aligns with the philosophy of India’s National Digital Health Mission (National Health Authority and Family Welfare, 2020). One solution could be to allow users to upload relevant documents or manually enter specific details they want the bot to access. This can empower users to control shared information while benefiting from a personalized experience. Further, when developing CataractBot in May 2023, we chose to use GPT-4, as it was the leading LLM (OpenAI et al., 2023). However, recent open-source models like LLaMA have shown comparable performance to GPT in medical applications (Li et al., 2025). Incorporating these models into the CataractBot system would enable local processing of patient data, helping address privacy concerns.

### v6

@@ Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we developed CataractBot [^1] Our code is publicly available at https://github.com/microsoft/BYOeB.. CataractBot answers cataract surgery related questions instantly using an LLM to query a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 49 patients and attendants, 4 doctors, and 2 patient coordinators, CataractBot demonstrated potential, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on expert-mediated LLM bots.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Interaction designccs: Applied computing Health care information systemsccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
To address patients’ information needs, particularly during surgical treatments, we propose a chatbot solution integrated with medical professionals to provide accessible, reliable and comprehensive responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of smartphones and instant messaging services such as WhatsApp, this chatbot serves as a 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified this experts-in-the-loop chatbot approach by developing CataractBot, which answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (Lewis et al., 2020) with an LLM over a custom knowledge base (Section 4.1.2) curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features to enhance adoption and sustained use. Capitalizing on the widespread adoption of smartphones, we chose a mobile-first design paradigm, deploying the chatbot entirely on the ubiquitous WhatsApp platform. To cater to diverse hospital visitors with varying literacy levels and technological proficiency, we made CataractBot multilingual (supporting five languages, including English, Hindi, and Kannada) and multimodal (accepting both speech and text inputs). We use an LLM to process complex and ill-formed queries, generating instant responses from the curated knowledge base. Crucially, CataractBot incorporates a system where the LLM’s responses are asynchronously verified by experts and corrected if needed. Ophthalmologists verify answers to medical queries and patient coordinators to logistical ones. These expert-provided edits are used to update the knowledge base, to minimize future expert intervention.
@@
Our main contributions are: (1) The design and development of a novel LLM-powered, experts-in-the-loop, multilingual, multimodal chatbot, created in collaboration with doctors and hospital staff to assist cataract surgery patients with their information needs. (2) The deployment of our bot in a real-world multi-stakeholder setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers insights specific to CataractBot and, more broadly, LLM-powered experts-in-the-loop chatbots.
@@ 2.2. LLM Chatbots in Healthcare
We chose cataract surgery–one of the world’s most common procedures (McGhee et al., 2020)–as the use case for developing and testing our system. This decision was motivated by our longstanding collaboration with Sankara Eye Hospital, an eye hospital in India. Also, cataract surgery involves well-defined protocols and common patient queries, making it an ideal starting point for testing the bot’s ability to deliver accurate, standardized information from a curated knowledge base. However, our solution is versatile, and we believe it can effectively support a range of health conditions, including mental health and chronic care. Given our specific focus on cataract surgery, we examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yılmaz and and (2025) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
@@ 3.2. Design Requirements
[…]
@@ 4. CataractBot System Design
Figure 1. CataractBot provides an initial response to the patient/attendant by querying the knowledge base. The doctor (or coordinator, if the question is logistical) verifies and corrects this response, and the patient/attendant is notified.
[…]
CataractBot is displayed with a split-screen view showing both patient/attendant (left) and doctor (right) perspectives. The workflow demonstrates: 1) A patient asks ”How many days after surgery can I wash my hair?”; 2) CataractBot initially responds that shampooing is allowed 3 days after surgery; 3) The doctor reviews this response, knowing it is intended for patient Jane Parvathi Doe (64/F/OD) who had cataract surgery on 2023-12-01, and marks it as incorrect; 4) The doctor provides the correct guidance, without having to worry about spelling or grammar, ”Btr avoid for 2 wks..”; 5) The patient is first notified that the bot’s answer was invalid and to wait for the doctor; and 6) Finally, the patient receives the doctor-verified corrected message stating ”Better to avoid washing your hair for 2 weeks after the cataract surgery.”
@@ 4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk (e.g., greetings or expressions of gratitude), or ‘other’, and provides a response in real-time. For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response. This custom knowledge base comprises of various cataract surgery and hospital-specific documents curated by hospital staff (doctors, patient coordinators, and members of Quality Control and the Patient Safety team). These documents, comprising approximately 30 pages in total, include the Consent Form, Standard Operating Procedures, FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. Some documents, like Post-Operative Guidelines, were included in patients’ discharge summaries. Others, such as SOPs and FAQs, existed in the hospital ecosystem but were not directly accessible to patients. The knowledge base also included verified question–answer pairs that were added over the course of the deployment. We provide details of this incremental updating process in Section 4.1.4. Grounded in this custom knowledge base, the bot-generated response includes a question mark icon [Uncaptioned image] as a ‘reaction’ indicating the unverified status (Figure 4E in Appendix A.1.2). In instances where the knowledge base lacks an answer, the bot responds with a template “I don’t know” response. For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
@@ 4.2. Implementation Details
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach of translating Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator (Microsoft, 2024). It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and knowledge base expert) get logged in a cloud storage for further analysis.
Figure 2. Overview of CataractBot System
[…]
This image shows a workflow diagram for a CataractBot. The process begins with an Information Seeker on the left who submits a Question, which goes through a Question Classifier. The classified question then queries a Knowledge Base, which outputs Data Chunks that feed into an LLM, which outputs an (unverified) response, shared with an Expert for verification or edits. The resulting output is sent back to the Information Seeker as a Verified Response with Related Questions. Knowledge Base Update Experts can use the Expert edits to provide Updates to the Knowledge Base.
@@ 5.1. Procedure for Patients and Attendants
[…]
@@ 6.1.1. Reasons for (Not) Using CataractBot
Figure 3. A. (Left) Messages sent per day pre- and post-surgery. B. (Top right) Messages sent before surgery, on the day of surgery, and after surgery. C. (Bottom right) Message input modality.
[…]
Figure A (left) displays message categories sent pre-surgery, during surgery, and post-surgery, with medical procedure questions dominating pre-surgery, medication and pain concerns prominent during surgery, and recovery questions increasing post-surgery. Figure B (top right) shows message frequency relative to surgery day (Day 0), revealing a sharp communication spike in the days immediately before and after surgery, with peak activity occurring on surgery day and decreasing gradually afterward. Figure C (bottom right) compares communication methods between patients and attendants, showing patients primarily use text messaging while attendants more frequently utilize a combination of text, audio, and taps.
@@ 7. Discussion
Below, we raise critical insights and open questions that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale across domains.
7.1. Scaling and Generalization
@@
Our open-sourced framework can be adapted beyond cataract surgery to support more complex, atypical, or long-term treatments–e.g., glaucoma surgery, cancer treatment, or pregnancy–where easy access to verified information would be valuable for patients, improving their health literacy and enabling better self-care. While conducting such studies, aligning the long-term research agenda with the operational priorities of busy hospitals could pose challenges, as cautioned by Agapie et al. (2024). It is crucial to ensure “that clinicians are fully aware of the motivations and methodologies of the (research) process, which is very different from a normal clinical situation” (Newell and Gregor, 2000). For instance, during our study, securing trust, recruiting hospital staff, and minimizing administrative delays required active endorsement from a senior doctor. This highlights the need to thoughtfully leverage existing authority structures in institutional settings. We also observed that positioning the technology as a long-term solution to reduce workload helped minimize clinicians’ resistance during the initial onboarding phase.
Beyond the medical domain, the applicability of our framework can be extended further, to fields such as law, finance and education. In these scenarios, end-users (e.g., defendants, taxpayers, or students) could receive synchronous responses from an LLM using a custom knowledge base, and subsequent asynchronous verified responses from experts (e.g., lawyers, accountants, or teachers). Expert-mediated bots can even be used to upskill less-trained workforce members. For instance, these bots could serve as valuable learning resources for community health workers, who typically have minimal medical training (Yadav et al., 2019).
7.2. Expert-in-the-loop Approaches
@@
Doctors are a scare resource in healthcare ecosystems. Our chatbot system can help scale their impact, by reducing the time spent answering routine questions at length during consultations and calls, thus freeing up time for additional patient care. However, as doctors had to verify responses during deployment, CataractBot also added to their workload. To minimize this, one approach is to deliver pre-verified answers to specific questions that the expert has repeatedly marked as correct. Given the potential for LLM hallucinations (Denecke et al., 2024), cached responses must be carefully matched to similar patient demographics and question phrasing before reuse. Moreover, the information seeker should retain agency within this process. If a pre-verified answer is deemed insufficient or irrelevant, they should be able to flag it to initiate the expert-in-the-loop workflow for a revised, verified response. Another approach is to use a second LLM for quality control while generating each response, as GPT-4’s evaluations of AI-generated responses align well with human expert assessments (Gumma et al., 2024). This maker-checker verification module would enable tracking of the bot’s response quality, especially as the knowledge base expands. The module could also evaluate the confidence level in the generated response, or the perceived urgency of the question, which could be used to prioritize sending nudges to experts for verification.
7.3. Multimodality and Personalization
@@
Our system faced challenges in accurately interpreting questions from voice messages due to limitations in transcription and translation technologies. The same inaccurate transcriptions were shared with doctors, leading to confusion. Such issues are less likely when deploying a voice-based input modality to a predominantly English-speaking population or in tier-1 language settings in the Global North, as seen in prior work (Yang et al., 2024). For future deployments of LLM-powered bots, particularly in low-resource languages, we propose the following enhancements: (1) implement a dictionary of common errors across languages to serve as a look-up resource, (2) provide the LLM with both the original query and its English translation to improve answer generation, and (3) display a transcription of any audio question back to the information seeker, allowing them to verify or clarify what the bot understood. These would serve as additional safeguards in cases where language technologies fail. Additionally, non-verbal cues in audio messages–such as signs of anxiety, impatience, or the presence of others–were lost in textual transcriptions. This hindered experts’ ability to deliver appropriate care (also noted in (Li et al., 2023)). To address this, we suggest making the original audio clips available to experts on demand.
Beyond audio, recent LLM models like GPT-4V offer visual capabilities that could enable such bots to process image-based inputs. This aligns with the recommendations of Tseng et al. (2023), suggesting that healthcare information can often be communicated more effectively through images and videos than through text alone. For example, a patient could upload a photo of a specific medicine to ask questions about it, instead of providing a written description. While this could improve the user experience for information seekers, it may increase the workload for experts. Further, prior research (Hao et al., 2024) cautions that although LLMs are proficient in text processing, their ability to handle other forms of media, such as image and video, remains less reliable.
@@
Moving on to personalization, our current CataractBot deployment integrates minimal sensitive data, such as the patient’s demographic details and date of surgery. Going forward, while tailored responses can boost adoption and utility by providing more relevant information, they also raise privacy concerns, potentially reducing usage among certain users. This issue may be more pronounced in the developing world, where there is more awareness around privacy. Additionally, experts may face increased workloads as they reference user-specific information for verification, as many personalized edits may not be suitable for inclusion in the generic knowledge base. Low (tech-)literate users might require assistance in making informed decisions about the use of their data (Okolo et al., 2021). A binary approach of either sharing all records or none may not address individual preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access. This aligns with the philosophy of India’s National Digital Health Mission (National Health Authority and Family Welfare, 2020). One solution could be to allow users to upload relevant documents or manually enter specific details they want the bot to access. This can empower users to control shared information while benefiting from a personalized experience. Further, when developing CataractBot in May 2023, we chose to use GPT-4, as it was the leading LLM (OpenAI et al., 2023). However, recent open-source models like LLaMA have shown comparable performance to GPT in medical applications (Li et al., 2025). Incorporating these models into the CataractBot system would enable local processing of patient data, helping address privacy concerns.